OpenStudy (anonymous):

If r is irrational, then 1/r is irrational prove

OpenStudy (anonymous):

use the contrapostive

myininaya (myininaya):

lets prove: if 1/r is rational, then r is rational

OpenStudy (anonymous):

what myininaya said

myininaya (myininaya):

what satellite said

OpenStudy (anonymous):

I see a waterloo math student. O_o

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

we should work together =P

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is mcw also one?

myininaya (myininaya):

do you still need or require more help?

OpenStudy (anonymous):

so how would u prove the contrapossitive of than

OpenStudy (anonymous):

+1

myininaya (myininaya):

if 1/r is rational then it could be written as a/b where a and b are integers (we are assuming r does not =0) if 1/r=a/b then r=b/a therefore r is rational

OpenStudy (anonymous):

contrapositive of if P then Q is if not Q then not P and they are equivalent. so you show that if 1/r is rational, then so is r. that proves it

OpenStudy (anonymous):

what myininaya said

myininaya (myininaya):

:)

OpenStudy (anonymous):

thx